I need help solving this expression...-6x-10=-96 I would appreciate any help as this is confusing the heck out of me!
Answer by mananth(16949)

-6x-10=-96
..
add +10 to both sides
-6x-10+10=-96+10
-6x=-86
divide both sides by -6
-6x/-6=-86/-6
..
x=86/6
x=43/3
...
m.ananth@hotmail.ca